# MindShift: Leveraging Large Language Models for Mental-States-Based Problematic Smartphone Use Intervention

### v1

@@ 2.3. Persuasion for User Behavior Change and Large Language Models
Persuasion is a psychological approach designed to influence attitudes, beliefs, or behaviors(Cialdini and Cialdini, 2007). Language is the most commonly used means of persuasion, and this approach leverages facts, emotional appeals, and so on in a message to achieve its goal. Its effectiveness has been shown in multiple fields, such as advertising to encourage consumers to buy a product(Bernritter et al., 2017), supporting mental health such as coping with stress (Morris et al., 2018; Orji and Moffatt, 2018), and managing physical health such as reducing snacking behavior (Kaptein et al., 2012). For smartphone use intervention, persuasion usually appears as reminders, such as leveraging the user’s usage time in a template format (Apple, 2021; Inc, 2023), or some thought-provoking statements (Xu et al., 2022b). Prior work has suggested that the personalization of persuasive content can enhance the effectiveness of intervention (Choe et al., 2017; Harrison et al., 2015; Kaptein et al., 2012). Some studies pointed out that the effectiveness of static, unchanging intervention methods would decrease over time, while dynamic intervention methods with diverse intervention timings or contents (even appearing randomly) can improve effectiveness (Kovacs et al., 2018).
The recent emergence of Large Language Models (LLMs), such as ChatGPT(OpenAI, 2022) and PaLM(Chowdhery et al., 2022), has made vast progress in the generation of personalized and diverse content. Recent studies have explored various health applications supported by LLMs, such as health information seeking (Yunxiang et al., 2023), mental health support (Lamichhane, 2023; Kumar et al., 2023), personal health coaching (Wei et al., 2023; Montagna et al., 2023), and health education (Montagna et al., 2023; Kung et al., 2023). Recent research in the HCI community has also explored LLMs’ potential in public health interventions (Jo et al., 2023). However, there is a gap in leveraging LLMs for a real-time, context-aware behavior intervention. In this work, we aim to leverage LLMs’ ability to dynamically generate personalized persuasion content to help users reduce their problematic phone usage.
2.4. Theoretical Background and Mental Models of Smartphone Use
Understanding what leads to problematic phone use is essential for the design of effective persuasion strategies. The Dual Systems Theory (Hofmann et al., 2009) has been used to explain the phone usage patterns (Pinder et al., 2018). This theory divides human cognitive activities into two types: System 1 (fast, intuitive, unconscious) and System 2 (slow, analytical, conscious). Problematic smartphone use is typically driven by System 1 (Hagger, 2016), as it mainly involves unconscious, rapid responses and is easy to be guided by instant gratification. Recent research suggests that two key factors contribute to the failure of users to act on their goals: (i) limited ability of System 2 control; and (ii) fluctuations caused by emotional states and fatigue (Lyngs et al., 2019). These findings suggest that mental states play a significant role in habitual smartphone usage(Pinder et al., 2018). While previous research has identified external contexts that may trigger the habitual use of smartphones (Tran et al., 2019), limited research has explored users’ mental states as a factor. In our work, we pinpoint the major mental states linked to habitual use and propose corresponding persuasion strategies for these mental states.
@@ 3. Mental States of Habitual Smartphone Use and Persuasion Strategies
To better understand the mental states of users’ smartphone use and guide the design of our intervention system, we first conducted a Wizard-of-Oz (WoZ) study, followed by a semi-structured interview study (Sec. 3.1). We summarize the results in Sec. 3.2. Based on our initial findings, we devise four persuasion strategies (Sec. 3.3).
@@ 3.1. Exploratory Wizard-of-Oz & Semi-structured Interview Studies
To identify particular smartphone usage behaviors requiring targeted interventions and further develop persuasive intervention content material, we first conducted a WoZ study in the wild, involving 12 end-users (6 females and 6 males, aged 18-28), followed by a semi-structured interview with another 10 users (5 females and 5 males, aged 18-29).
@@ 3.1.1. Wizard-of-Oz Study
At the beginning of the study, participants were told that it was an automatic chatbot instead of a human to reduce the observation effect (Kazdin, 1982). In reality, when participants opened a blacklisted app, a human experimenter would receive an email notification. Based on participants’ smartphone usage data, the experimenter designed and delivered persuasive messages aiming to reduce smartphone usage via the chatbot. Each evening, researchers carried out a brief daily interview to probe (1) why participants use smartphones, (2) factors affecting their acceptance of interventions, and (3) evaluation of the effectiveness of the persuasive messages. Our WoZ study lasted five days for each participant. At the study’s conclusion, we informed participants that the chatbot was operated by a human experimenter at the back end.
@@ 3.1.2. Semi-structured Interview Study
We also carried out a semi-structured interview study with a different participant group. During the semi-structured interview, we asked participants to recollect instances of problematic smartphone use. Next, we asked them to discuss their mental states during those instances. We followed the participants’ lead during the interview. All interviews were audio recorded and transcribed. Three researchers independently examined the transcriptions and then collectively coded them through multiple iterations until they reached a consensus.
@@ 3.2. Main Takeaways about Problematic Smartphone Use
First, interventions for problematic smartphone use should target habitual usage. Our WoZ study delineates two primary types of smartphone usage: instrumental and habitual, which corroborates the classification of prior research  (Lukoff et al., 2018). We find that only habitual smartphone use warrants intervention. Providing interventions during instrumental use often leads to user dissatisfaction. Notably, relaxation emerges as a crucial form of instrumental use, where participants deliberately use their phones to unwind or reward themselves after intense work or study. Intervening at such times is considered intrusive and inappropriate. This classification aligns with the Dual Systems Theory. In smartphone interactions, instrumental use relies on conscious decision-making (System 2), while habitual use is more instinctive (System 1). Hence, interventions should primarily target habitual use, a conclusion also supported by earlier studies  (Lukoff et al., 2018).
@@
Second, the effectiveness of interventions depends on users’ mental states, personal goals, and contextual information. This is consistent with the literature, suggesting that a shift away from System 2 is due to emotional fluctuations and the absence of defined goals and intentions (Lyngs et al., 2019). We experimented with different persuasive message content during our WoZ study. We find that when we incorporate users’ mental states factor, which is inferred by their physical contexts and app usage patterns, into persuasive message content, participants have an increased acceptance rate of the intervention. In addition, we find emphasizing users’ personal goals can enhance the interventions’ effectiveness. For instance, sending persuasive messages to participants such as, “When you find yourself with idle time, consider engaging in meaningful activities such as reading, writing, or drawing” proves effective when users are in an idle state and aimed for self-improvement. Our findings are supported by prior studies linking habitual smartphone use to specific mental states  (Anderson and Wood, 2021; Bayer and LaRose, 2018).
@@
Third, semi-structured interviews reveal three primary mental states connected to habitual smartphone use: boredom, stress, and inertia.
@@
  • Boredom is an affective state characterized by low arousal and dissatisfaction due to insufficient stimulation  (Fisherl, 1993; Mikulas and Vodanovich, 1993). The WoZ and interview studies identify common scenarios leading to boredom: (1) when the task at hand is too simple, lacking a balance between skill and challenge, (2) lack of interest in the current activity, and (3) devoid of any engaging activities during idle moments, which can trigger cognitive biases like the fear of missing out and promote habitual use  (Bozaci, 2020; Smith, 2015; Wolniewicz et al., 2020).
@@
  • Stress refers to cognitive and behavioral reactions to unpredictable and unmanageable stimuli (Koolhaas et al., 2011). Participants frequently use smartphones because they experience (1) heightened anxiety when work demands exceed their abilities and (2) uncertainty about whether something would have a positive outcome. This aligns with increasing evidence links smartphone use to perceived stress (Chiu, 2014; Samaha and Hawi, 2016; Vahedi and Saiphoo, 2018).
@@
  • Inertia, in our context, refers to a psychological resistance that makes users reluctant to change their current activity state. It is similar to the idea illustrated by literature such as emotional inertia (Kuppens et al., 2010) or decision-making inertia (Alós-Ferrer et al., 2016). Our research shows that participants commonly use their phones habitually to avoid changing into a new activity state from an idle state (e.g., checking the phone before starting to work). Unlike stress or boredom, inertia does not elicit overt negative emotions but impedes the shift to the next task.
Moreover, we find that users’ mental states are either triggered by specific activities or have no specific reason. Therefore, we categorize scenarios based on whether users are engaged in specific activities (“engaging in activities”, “not engaging in activities”). The first category (“engaging in activities”) denotes mental states triggered by ongoing or upcoming activities. For instance, activities that lack challenge and engagement, or activities that involve repetitive actions might induce boredom. In contrast, the second category (“not engaging in activities”) has less clear causes, often linked to broader life situations. For example, a lack of enthusiasm for life may cause boredom, while challenges adapting to setbacks and uncertainties can cause stress. Hence, combining the three mental state factors with whether or not engaging in activities factor, we categorize habitual phone use into six distinct scenarios (see the gray part of Figure 2).
Figure 2. Summary of Users’ Mental States and Corresponding Persuasion Strategies Based on WoZ and Interview Studies. In the 3 x 2 table (three mental states: boredom vs. stress vs. inertia, and two activities: engage in vs. not engage in), each color denotes the persuasion strategy applicable in this scenario.
@@ 3.3. Persuasion Strategies Design
Based on the Dual Systems Theory and the mental states we summarized in Sec 3.2, we propose four distinct persuasion strategies: Understanding, Comforting, Evoking, and Scaffolding Habits (see the colorful part in Figure 2).
Our strategy design is inspired by the theory of Existence, Relatedness, and Growth (ERG). Developing from Maslow’s Hierarchy of Needs, ERG theory further summarizes human motivation into three levels: (i) the physiological and safety basic needs for existence, (ii) the social needs for feeling related and accepted, and (iii) the need to grow and self-actualize. The theory has been applied to workplaces to increase productivity and job satisfaction (Arnolds and Boshoff, 2002; Caulton, 2012).
Maintaining System 2 relies on cultivating motivation for meaningful activities. Thus, we map strategies to the relatedness level and growth level of the ERG theory in Figure 3. At the level of relatedness, Understanding and Comforting aim to empathize with users’ emotions, offering support and empowering users to manage System 2. At the growth level, Evoking reminds users of their personal development goals, and Scaffolding Habits guides them in replacing habitual smartphone use with activities conducive to self-fulfillment, thereby turning awareness into action.

### v2

@@ 2.4. Persuasion for User Behavior Change and Large Language Models
Persuasion is a psychological approach designed to influence attitudes, beliefs, or behaviors (Cialdini and Cialdini, 2007). Language is the most common means of persuasion (Simons, 1976; Rapp, 2002), and leverages facts, emotional appeals, and so on to achieve its goal. Its effectiveness has been shown in multiple fields, such as advertising to encourage consumers to buy a product (Bernritter et al., 2017), supporting mental health such as coping with stress (Morris et al., 2018; Orji and Moffatt, 2018), and managing physical health such as reducing snacking behavior (Kaptein et al., 2012). For smartphone use intervention, persuasion usually appears as reminders, such as leveraging the user’s usage time in a template format (Apple, 2021; Inc, 2023), or some thought-provoking statements (Xu et al., 2022b). Prior work has suggested that personalizing content can enhance persuasion effectiveness such as reducing snacks  (Kaptein et al., 2015, 2012). Also, varying the timing and content of interventions, sometimes even randomly, can improve effectiveness. In contrast, static interventions tend to lose influence over time (Kovacs et al., 2018).
The advent of Large Language Models (LLMs), like ChatGPT(OpenAI, 2022) and PaLM(Chowdhery et al., 2022), has made vast progress in personalized and diverse content generation. Recent studies have explored various health applications supported by LLMs, such as health information seeking (Yunxiang et al., 2023; Mahmood et al., 2023; Yang et al., 2023), mental health support (Lamichhane, 2023; Kumar et al., 2023; Xu et al., 2023), personal health coaching (Wei et al., 2023; Montagna et al., 2023), health education (Kung et al., 2023), and public health interventions (Jo et al., 2023). These applications showcase LLMs’ capabilities in knowledge delivery and emotional support. Compared to them, our study further explores LLMs for just-in-time behavior change and intervention, beyond information presentation.
@@ 3. Mental States of Habitual Smartphone Use and Persuasion Strategies
To comprehend the mental states of users’ smartphone use and guide our intervention system design, we initiated a Wizard-of-Oz (WoZ) study, followed by a semi-structured interview study (Sec. 3.1). We summarized the main takeaways in Sec. 3.2. Based on theories and our findings, we devised four persuasion strategies and their implementation under different mental states (Sec. 3.3).
@@ 3.1. Exploratory Wizard-of-Oz & Semi-structured Interview Studies
To identify particular smartphone usage behaviors requiring targeted interventions, we first recruited 12 end-users (6 females and 6 males, aged 18-28) and conducted a 5-day WoZ study in the wild. The findings suggested ideas for persuasion content design. For deeper insights into participants’ mental states and concrete intervention design materials, we recruited another group of 10 users (5 females and 5 males, aged 18-29) and conducted a semi-structured interview study [^1] Since the two studies are close in time, we choose a separate set of participants for semi-structured interviews to avoid the impact of intervention in the WoZ study.. Both studies are approved by the institution’s IRB. Our studies focused on young adults who have been reported to have the most severe problematic smartphone use issues (Mitchell and Hussain, 2018; Lee et al., 2014). However, we do recognize that our sampling could limit the generalizability of our findings. We discuss this as a limitation in Sec. 8.
@@ 3.1.1. Wizard-of-Oz Study
To reduce the observation effect, participants were told that it was an automatic chatbot instead of a human (Kazdin, 1982). In reality, when participants opened a blacklisted app, a human experimenter would receive an email notification. Based on smartphone usage duration and frequency (see the detection method in Sec. 5), the experimenter designed and delivered persuasive messages. Inspired by existing literature on persuasion design (Ham and Langrial, 2020; Albarracín et al., 2018; Block and Keller, 1997), our messages fell into 4 types (see examples in Table 4 in Appendix): (1) usage notice, telling participants their usage data such as the accumulated usage time today and time since last use, (2) practical guidance, asking participants’ goals today and suggesting tasks instead of smartphone use, (3) encouragement, praising and cheering participants to keep smartphones away, and (4) deterrent, alarming participants the consequences of using smartphones such as task delay and admonishing them to stop.
Every evening, researchers conducted a brief 15-minute online interview in person with each participant, structured around four questions: (1) What was your overall experience of using the chatbot? How did it change your smartphone usage? (2) Why were you using your phone at a particular time? (3) What were your reactions to the persuasive message, and why? (4) How did you like the persuasive message? How can it be improved? At the study’s conclusion, we informed participants that the chatbot was actually operated by a human experimenter at the back end.
All persuasive messages and participants’ responses, along with their sending times, are documented. Daily interviews were audio-recorded and transcribed. The recorded data and transcriptions were independently reviewed by three researchers, who coded them based on two main themes: types of smartphone use for question (2) and factors influencing persuasion effectiveness for questions (1), (3), and (4). Subsequently, the researchers convened to discuss the codes until a consensus was reached. Following that, a thorough review of all transcriptions was conducted to ensure the accuracy of the coding.
@@ 3.1.2. Semi-structured Interview Study
Our WoZ study provided insights into participants’ problematic use behavior and reactions towards persuasion. To obtain a deeper understanding of the user’s mental states during phone use, we conducted a semi-structured interview study with another participant group. We asked participants to recollect instances of problematic smartphone use. Our interview started with the question: “When would you want to use an intervention app to limit your smartphone use?” We then sought details about the scenario (e.g., time, place, and concurrent activities) and user behaviors and reactions (e.g., usage duration, feelings, and reflections). Next, we asked participants to share their mental states during those instances. We asked questions: “Why do you use your phone even though you think you should not? What’s your mental state behind these reasons?” We followed the participants’ lead during the interview.
All interviews were audio recorded and transcribed. Three researchers independently examined the transcriptions and coded the mental states in different smartphone use cases and contexts. Then they met and discussed the codes until reaching a consensus. To ensure coding accuracy, they went through all transcriptions one more time.
@@ 3.2. Main Takeaways about Problematic Smartphone Use
We summarized our main takeaways from the WoZ study and the interview study below. Table 5 in the Appendix summarizes our findings with participants’ quotes. Table 1 summarizes the findings about mental states in Takeaway \⃝raisebox{-0.9pt}{3} and Takeaway \⃝raisebox{-0.9pt}{4}.
Table 1. Summary of Users’ Mental States behind Habitual Smartphone Use.
Boredom
Engaging in Activities Not Engaging in Activities
Users find current tasks boring, lack interest, and struggle to concentrate. This could be due to the task lacking challenges, not being sufficiently engaging, having high repetition, and not aligning with the user’s genuine interests and desires.
Users feel bored with daily living in general, lack passion, and have no enthusiasm for engaging in activities. This might be because they have nothing to do, don’t know how to pass the time, lack excitement in life, and feel that living is meaningless.
Stress
Engaging in Activities Not Engaging in Activities
Users feel stressed about current tasks due to challenges posed by the environment, which deplete their resources and result in feelings of tension and unhappiness. This might be due to the abundance, difficulty, and urgency of tasks, causing users to feel anxious, fatigued, and lacking confidence in their abilities, leading to a pessimistic view of the outcomes.
Users feel stressed in the face of daily living and challenges from the environment that deplete their resources, making them feel tense and unhappy. This might be due to setbacks and unexpected events in life that users struggle to adapt to, leading to a pessimistic outlook on the future.
Inertia
Engaging in Activities Not Engaging in Activities
Users find it difficult to transition from their current state to start the next activity, but without explicit negative emotions. This might be due to procrastination has become a habit, and there’s insufficient motivation for the next activity.
Users indulge in idle inertia, but without specific negative emotions. This might be due to idling around has become a habit, and there’s no motivation to organize new activities.
Others
Beyond the scope of the current paper.
The table is organized into seven rows and two columns, including headers. The first, third, and fifth rows combine the two columns with the headers ”Boredom”, ”Stress”, and ”Inertia” respectively. In the second, fourth, and sixth rows, the left column contains concrete descriptions of boredom, stress, and inertia under the condition of ”engaging in activities”; the right column contains concrete descriptions of boredom, stress, and inertia under the condition of ”not engaging in activities”. The seventh row combines the two columns with the header ”Others”, referring to mental states ”beyond the scope of the current paper”.
Table 1. Summary of Users’ Mental States behind Habitual Smartphone Use.
@@
Takeaway \⃝raisebox{-0.9pt}{1} Interventions for problematic smartphone use should target habitual usage. Our WoZ study delineated two primary types of smartphone usage: instrumental and habitual, consistent with prior research categorization  (Lukoff et al., 2018). We found that only habitual smartphone use warrants intervention. Interventions during instrumental use often led to user dissatisfaction. Notably, relaxation emerged as a crucial form of instrumental use, where participants deliberately used their phones to unwind or reward themselves after intense work or study. Intervening at such times was considered intrusive and inappropriate. The finding aligns with the Dual Systems Theory. In smartphone interactions, instrumental use relies on conscious decision-making (System 2), while habitual use is more instinctive (System 1). Hence, interventions should primarily target habitual use, which is also supported by earlier studies  (Lukoff et al., 2018).
@@
Takeaway \⃝raisebox{-0.9pt}{2} The effectiveness of interventions depends on the alignment with users’ mental states, personal goals, and contextual information. This is consistent with the literature, suggesting that a shift away from System 2 is due to emotional fluctuations and the absence of defined goals and intentions (Lyngs et al., 2019). We experimented with different persuasive message content during our WoZ study. We found that when we incorporated users’ mental states as a factor, which was inferred by their physical contexts and app usage patterns, into generating persuasive message content, participants were more willing to accept the intervention. Furthermore, highlighting users’ personal goals enhanced intervention effectiveness. For instance, sending messages like, “When you find yourself with idle time, consider engaging in meaningful activities such as reading, writing, or drawing” proved effective when users were in an idle state and had a goal for self-improvement. Our findings are supported by prior studies linking habitual smartphone use to specific mental states  (Anderson and Wood, 2021; Bayer and LaRose, 2018).
@@
Takeaway \⃝raisebox{-0.9pt}{3} Semi-structured interviews revealed three primary mental states connected to habitual smartphone use: boredom, stress, and inertia.
@@
  • Boredom is an affective state characterized by low arousal and dissatisfaction due to insufficient stimulation  (Fisherl, 1993; Mikulas and Vodanovich, 1993). The WoZ and interview studies identified common scenarios leading to boredom: (1) when the task at hand is too simple, lacking a balance between skill and challenge, such as ”doing simple assignments light on cognitive engagement” (S3) [^2] This is the serial number of the participant in the semi-structured interview study. , (2) lack of interest in the current activity, such as ”completing assignments is to relieve a burden, instead of reaching achievement” (S6), and (3) devoid of any engaging activities during idle moments, such as ”after returning to home” when is ”not yet time to sleep” (S8).
@@
  • Stress refers to cognitive and behavioral reactions to unpredictable and unmanageable stimuli (Koolhaas et al., 2011). Participants frequently use smartphones because they experience (1) heightened anxiety when work demands exceed their abilities, such as ”having a challenging bug to locate when doing programming assignment” (S3) or ”work not progressing well” (S9) and (2) uncertainty about whether something would have a positive outcome, such as ”not sure if can get a job offer” (S10). This aligns with increasing evidence that links smartphone use to perceived stress (Chiu, 2014; Samaha and Hawi, 2016; Vahedi and Saiphoo, 2018).
@@
  • Inertia, in our context, refers to a psychological resistance that makes users reluctant to change their current activity state. It is similar to the idea illustrated by literature such as emotional inertia (Kuppens et al., 2010) or decision-making inertia (Alós-Ferrer et al., 2016). Participants stated they commonly used their phones habitually to avoid changing into a new activity state from an idle state, ”checking the phone before starting to do assignment” (S1) or ”shifting from a relaxed state to a focused state” (S2). Unlike stress or boredom, inertia does not elicit overt negative emotions but impedes the shift to the next task.
Takeaway \⃝raisebox{-0.9pt}{4} Engaging vs. Not Engaging in Activities (Table 1). We further noticed two nuanced categories within mental states. The first category (“engaging in activities”) denotes that users have activities to complete while habitually using smartphones. Participants either got distracted from the ongoing boring or stressful activities (e.g., ”I find myself instinctively reaching for my phone in search of mental stimulation when doing simple assignments light on cognitive engagement” (S3)) or procrastinated to face the upcoming activities (e.g., ”I was reluctant to start handling this challenging work that I scrolled my phone screen anxiously” (S2)). In contrast, the second category (“not engaging in activities”) means users have no schedule or don’t know what to do while using smartphones habitually (e.g., ”After getting off work and returning home, I collapse on the sofa and binge-watch Tiktok for one to two hours” (S9)). Differentiating activity engagement states and combining with the three identified mental states lead to six granular categories of habitual smartphone use, making the persuasion strategies design more situated to users’ scenarios. For users engaging in activities, the persuasion not only aims to stop them from smartphone use but also to encourage them to either continue or initiate their activities.
@@ 3.3. Persuasion Strategies Design
Based on the takeaways and inspired by the Dual Systems Theory and the Existence, Relatedness, and Growth (ERG) theory, we proposed four distinct persuasion strategies: Understanding, Comforting, Evoking, and Scaffolding Habits. Developing from Maslow’s Hierarchy of Needs, the ERG theory further summarizes human motivation into three levels: (i) the physiological and safety basic needs for existence, (ii) the social needs for feeling related and accepted, and (iii) the need to grow and self-actualize. The theory has been applied to workplaces to increase productivity and job satisfaction (Arnolds and Boshoff, 2002; Caulton, 2012).
Some of our takeaways align with these theories. For example, Takeaway \⃝raisebox{-0.9pt}{1} aligns with the Dual Systems Theory and suggests that to avoid habitual smartphone use out of instinctive System 1, it is necessary to cultivate enough motivation to maintain conscious but difficult System 2. Moreover, Takeaway \⃝raisebox{-0.9pt}{2} shows that persuasive messages relieving mental states and reminding personal goals are effective, which is consistent with the human motivation of relatedness and growth outlined in the ERG theory.
Accordingly, we map strategies to the relatedness level and growth level of the ERG theory to arouse users’ motivation for System 2 (Figure 2). The existence level concerning physiological and safety needs is not included in our theoretical framework. At the level of relatedness, Understanding and Comforting aim to empathize with users’ emotions, offering support and empowering users to manage System 2. At the growth level, Evoking reminds users of their personal development goals, and Scaffolding Habits guides them in replacing habitual smartphone use with activities conducive to self-fulfillment, thereby turning awareness into action. Then, we map 4 strategies to 3×2 mental states in Figure 3.
@@
This is a picture of a balance scale. On the left side of the scale, there is a square labeled ”System 1,” and on the right side, there is a square labeled ”System 2.” These two squares are enclosed by dashed lines and labeled ”Dual Systems Theory,” indicating that we are applying the Dual Systems Theory. On the square labeled ”System 2,” there are also squares stacked in sequence with labels ”Relatedness Needs” and ”Growth Needs,” and these two squares are enclosed by dashed lines and labeled ”ERG Theory,” where the letter ”E” is colored gray, lighter than the other black letters, indicating that we are only using the second layer (relatedness) and the third layer (growth) of the ERG Theory. The right side of the scale sinks lower than the left side, and the square on System 2 is labeled ”Dominate.” This is meant to convey that after using the ERG Theory to stimulate user Related Needs and Growth Needs, System 2 can gain an advantage in competition with System 1. The square labeled ”Related Needs” has arrows pointing to two squares labeled ”Understanding” and ”Comforting” respectively. The square labeled ”Understanding” has an arrow pointing to the square labeled ”Comforting,” indicating that the Comforting intervention strategy goes a step further than the Understanding intervention strategy. The square labeled ”Growth Needs” points to two squares labeled ”Evoking” and ”Scaffolding Habits” respectively. The square labeled ”Scaffolding Habits” has an arrow pointing to the square labeled ”Evoking,” indicating that the Scaffolding Habits intervention strategy goes a step further than the Evoking intervention strategy.
Figure 3. Summary of Persuasion Strategies under Different Mental States. In the table, each color denotes the corresponding persuasion strategy applicable to this scenario.
This is a 3 x 2 table, with three mental states (boredom, stress, inertia) and two activity states (engaging in activities, not engaging in activities). After the ”three mental states” column, the fourth column is an ”Others” column, referring to mental states ”beyond the scope of the current paper”. At the bottom of the table, there are four squares labeled ”Understanding,” ”Comforting,” ”Evoking,” and ”Scaffolding Habits.” In the ”boredom × engaging in activities” cell, there are four rows of concrete descriptions of ”Understanding,” ”Comforting,” ”Evoking,” and ”Scaffolding Habits.” In the ”boredom × not engaging in activities” cell, there are three rows of concrete descriptions of ”Understanding,” ”Comforting,” and ”Scaffolding Habits.” In the ”stress × engaging in activities” cell, there are four rows of concrete descriptions of ”Understanding,” ”Comforting,” ”Evoking,” and ”Scaffolding Habits.” In the ”stress × not engaging in activities” cell, there are three rows of concrete descripyions of ”Understanding,” ”Comforting,” and ”Scaffolding Habits.” In the ”inertia × engaging in activities” cell, there are three rows of concrete descriptions of ”Understanding,” ”Evoking,” and ”Scaffolding Habits.” In the ”inertia × not engaging in activities” cell, there are two rows of concrete descriptions of ”Understanding,” and ”Scaffolding Habits.”
Figure 3. Summary of Persuasion Strategies under Different Mental States. In the table, each color denotes the corresponding persuasion strategy applicable to this scenario.